# Visual Document Understanding and Question Answering: A Multi-Agent Collaboration Framework with Test-Time Scaling

### v1

@@ Introduction
Since that scaling test-time computation can activate the long-context understanding and reasoning capabilities of small-parameter VLMs, dynamically allocating additional computing resources and enabling multiple paths in the test time, we propose an agent-wise hybrid scaling strategy during inference. Existing test-time scaling methods fall into four main categories: parallel scaling, sequential scaling, hybrid scaling, and internal scaling. However, these strategies are originally designed for single models or agents, overlooking the different division of labor among agents and achieving suboptimal performance when applied to multi-agent systems. Accordingly, we propose an agent-wise hybrid scaling strategy tailored to multi-agent architectures, which significantly enhances both agent-specific capabilities and collaborative performance.
@@
For the first three agents with various functions, we customize different test-time scaling strategies to their unique characteristics: (1) Planning Agent: In our design of $\mathcal{A}_{plan}$, it naturally provides relevant sample plans as references for formulating multiple execution plans for the original task. We therefore prompt the generation of $N_{p}$ relevant plans independently, yielding $N_{p}$ parallel paths per query. It establishes a foundation for scaling subsequent agents and increases the likelihood that at least one path will produce the correct answer. (2) Execution Agent: Given that the execution process is divided by step in $\mathcal{A}_{exe}$, we treat each step as an evaluation node. For each node, the agent produces $\mathit{N_{e}}$ candidate executions, which are scored using a pretrained reward model. The top-scoring candidate is selected as the base node for subsequent steps, with all others rejected. (3) Judgment Agent: To judge the correctness of the execution plans and processes, $\mathcal{A}_{judg}$ requires logical analysis and reasoning to comprehensively and accurately detect mistakes. We therefore adopt the budget forcing scaling method (Muennighoff et al. 2025) for this agent, which enforces a minimum number of thinking tokens. When token usage falls below the budget, the agent is encouraged to generate additional thinking tokens, thereby promoting accurate judgments. For the answer agent, whose key function is to summarize prior information and generate the final answer, test-time scaling yields only limited improvements, so we do not apply scaling to it.
Document Non-Document Average
@@
Average Images Per Question 1.0 5.7 20.0 47.5 1.0 1.2 1.0 1.0 1.0 1.0 0.5 1.0 1.0 1.0 1.0 5.7
Average Question Length 8.3 8.7 9.0 16.4 10.6 11.5 7.2 20.4 10.5 20.3 12.1 11.1 15.6 42.3 35.7 16.0
Average Answer Length 2.1 3.4 2.4 2.8 9.5 1.6 1.2 2.9 4.7 8.5 4.4 1.2 1.2 2.2 1.4 3.3
@@
MACT 96.6 72.5 85.3 43.7 92.0 89.4 91.9 85.2 74.0 57.2 78.5 77.7 81.2 41.8 54.7 74.8
-Qwen2.5-VL-Series-24B +2.0 +10.2 +9.1 +21.3 +5.2 +6.6 +4.4 +17.0 +13.6 +8.0 +4.5 +9.3 +13.4 +16.3 +13.9 +10.3
MACT 94.4 70.8 83.9 47.4 93.8 88.6 91.4 87.2 71.6 62.7 79.2 76.1 85.4 60.1 65.3 77.2
-MiMo-VL-Series-28B +1.5 +8.5 +9.7 +19.9 +6.5 +5.9 +5.3 +15.2 +14.1 +10.2 +4.1 +8.2 +14.7 +13.2 +12.0 +9.9
MACT 96.1 69.8 81.3 46.5 91.0 87.7 90.6 85.0 75.3 61.0 81.3 80.1 83.8 45.8 54.0 75.3
-InternVL3-Series-28B +5.0 +9.7 +8.4 +17.9 +6.6 +9.2 +5.2 +17.9 +15.6 +8.6 +4.9 +13.4 +18.3 +17.0 +15.1 +11.5
Table 1: DVQA, SVQA, MMLong, InfQA, VisMRC, TabVQA, TabBen, SciQA, RealQA, MVista, MVision, and MVerse denote DocVQA, SlideVQA, MMLongBench-Doc, InfographicVQA, VisualMRC, TableVQA-Bench, TableBench, ScienceQA, RealWorldQA, MathVista, Math-Vision, and MathVerse benchmarks, respectively. The values with green, blue, and gray colors denote the best, the second best, and the third best values. Results with - indicate exceeding the maximum context token limit or not being equipped with a specific ability. $\triangle,\Box,\Diamond$ correspond to the base models of MACT for comparisons.
Experimental Settings
Training Pipeline
@@
Since the functions and optimization goals of each agent are independent and diverse, we design a two-stage SFT and RL pipeline for MACT. To understand visual inputs and better accomplish their labors, we choose VLMs for the $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, while LLMs for the $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$. For all the four agents, we initiate with pretrained models, and we select three groups of small-parameter base models for different variants of MACT: (1) Qwen2.5-VL Series Based (Bai et al. 2025; Yang et al. 2025): Qwen2.5-VL-7B-Instruct and Qwen2.5-7B/3B-Instruct; (2) MiMo-VL Series Based (Xiaomi 2025; Xia et al. 2025): MiMo-VL-7B-SFT and MiMo-7B-SFT; (3) InternVL3 Series Based (Zhu et al. 2025): InternVL3-9B/8B/2B-Instruct. In the first SFT stage, we initially train a tuned 11B/7B/7B VLM on the selected document-based or non-document-based datasets, mixing data with or without CoT. It aims to enhance their visual understanding and reasoning abilities, providing more robust models for future multi-agent collaboration. Next, we fine-tune an 8B/7B/7B LLM on judgment labels generated via GPT-4o (Hurst et al. 2024) and rule-based verifications. Finally, we fine-tune another 3B/3B/7B LLM on the outputs from preceding agents and ground-truths to enhance its summary ability. The fine-tuned VLM are applied as $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$ respectively in the subsequent stage, while the two LLMs function as the $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$. In the second RL stage, we generate reward signals based on pretrained reward models and optimize our model via GRPO (Shao et al. 2024). Process reward model VisualPRM (Wang et al. 2025c) is utilized for the step-by-step reward signals of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, while Skywork-VL-Reward (Wang et al. 2025d) is used for the rewards generation of $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$. More training details are in the Appendix.
Datasets
@@
To train our model and comprehensively, objectively assess its visual capabilities, we selected 15 datasets spanning four real-world document types and two non-document types. The four document types are comprised of: (1) Text-Based: DocVQA (Mathew, Karatzas, and Jawahar 2021), DUDE (Van Landeghem et al. 2023), SlideVQA (Tanaka et al. 2023), MMLongBench-Doc (Ma et al. 2025); (2) Webpage-Based: VisualMRC (Tanaka, Nishida, and Yoshida 2021), InfographicVQA (Mathew et al. 2022); (3) Chart-Based: ChartQA (Masry et al. 2022), CharXiv (Wang et al. 2024d); (4) Table-Based: TableVQA-Bench (Kim, Yim, and Song 2024), and TableBench (Wu et al. 2025). To retain generalization and enhance reasoning, two non-document types are involved: (1) General: ScienceQA (Saikh et al. 2022), RealWorldQA (xAI 2024); (2) Mathematical: MathVista (Lu et al. 2024a), Math-Vision (Wang et al. 2024b), and MathVerse (Zhang et al. 2024c). We adhere to the original training and testing splits and use the test splits as evaluation benchmarks. Each instance in these datasets consists of a visual input and a question (except ScienceQA), spanning difficulty levels from easy to hard and encompassing various question types. More details are provided in the Appendix.
Evaluations
@@
Following (Li et al. 2025b), we use the reported initial results of reproducible models where configurations align with ours. For other models, we utilize LMMs-Eval (Zhang et al. 2024b) for fair comparisons on both natively supported and our registered benchmarks. For most benchmarks, we employ GPT-4o (Hurst et al. 2024) as a judge model to evaluate the correctness of the generated answers, which is achieved in LMMs-Eval. For the remaining benchmarks, we utilize their original evaluation metrics, e.g., ANLS and F1. More details can be seen in the Appendix.
Results and Discussions
Main Results
@@
We evaluate MACT across 15 benchmarks: 10 document-based benchmarks to assess its document understanding performance; 2 general benchmarks to gauge generalization ability, mitigating overfitting to specific scenarios; and 3 mathematical benchmarks to evaluate complex reasoning capabilities. For comparison, we select state-of-the-art (SOTA) methods—encompassing both generalist models and document-specific specialists—and categorize them by parameter size.
@@
As shown in Table 1, the MACT-MiMo-VL-Series-28B variant delivers the best average performance, followed by MACT-InternVL3-Series-28B and MACT-Qwen2.5-VL-Series-24B, which rank second and third, respectively. Despite having fewer than 30B parameters, MACT models outperform all comparative methods with under 100B parameters, as well as closed-source models. Notably, MACT-MiMo-VL-Series-28B achieves significant improvements of 5.6% and 5.9% over the top-performing open-source and closed-source models in terms of the average scores. Additionally, the three variants top 13 of the 15 benchmarks, with MACT-MiMo-VL-Series-28B leading on seven. Notably, in MMLongBench-Doc, which features the longest visual context, and across the three mathematical reasoning benchmarks, MACT-MiMo-VL-Series-28B outperforms the second-highest scorer by 7.1%, 10.6%, 5.9%, and 8.7%, respectively. These results highlight the superior overall performance of MACT, as well as its strong capabilities in long-context understanding and reasoning.
@@
Radar charts in Figure 1 compare three variants of MACT with their base models and larger models from the same series. As shown, MACT models significantly outperform their base models by 10.3%, 9.9%, and 11.5% on average across the 15 benchmarks. Compared to their corresponding larger-parameter counterparts, i.e., Qwen2.5-VL-72B-Instruct and InternVL3-78B-Instruct, our MACT-Qwen2.5-VL-Series-24B and MACT-InternVL3-Series-28B achieve average performance gains of at least 6.6% and 3.7%, respectively, with more pronounced improvements observed on long-context and reasoning tasks.
Additional Quantitative Analysis
@@
To further demonstrate the effectiveness of our methodologies, we perform ablation experiments and analyses on the proposed multi-agent collaboration, reward modeling, and test-time scaling strategies. Experiments here are conducted using MACT-Qwen2.5-VL-Series-24B. Specifically, we select three lowest-scoring benchmarks, which are characterized by long visual contexts and reasoning demands, and the average results across all benchmarks to assess their impact on both specific abilities and overall performance.
Ablation Studies
MMLong TabBen MVision Average
w/o Multi-Agent Collaboration 32.5 $\downarrow$11.2 50.8 $\downarrow$  6.4 32.4 $\downarrow$  9.4 66.2 $\downarrow$  8.6
w/o Multi-Agent Collaboration 24.7 $\downarrow$19.0 44.4 $\downarrow$12.8 26.0 $\downarrow$15.8 58.6 $\downarrow$16.2
w/o Mixed Reward Modeling 38.3 $\downarrow$  5.4 54.2 $\downarrow$  3.0 36.7 $\downarrow$  5.1 71.4 $\downarrow$  3.4
w/o Agent-Wise Hybrid Scaling 34.9 $\downarrow$  8.8 50.8 $\downarrow$  6.4 34.5 $\downarrow$  7.3 71.1 $\downarrow$  3.7
@@
Table 2: Results of ablations. $\dagger$ denotes a single agent that directly outputs answers, while $\ddagger$ indicates a single agent that incorporates all prompts as a workflow.
$\mathcal{A}_{plan}$ + $\mathcal{A}_{exe}$ $\mathcal{A}_{judg}$ $\mathcal{A}_{ans}$ MMLong TabBen MVision Average
@@
Table 3: Results of different combinations of agents.
The ablation settings are as follows: (1) w/o Multi-Agent Collaboration: Use a single agent to directly execute tasks and output answers, with our proposed scaling and reward strategies retained. (2) w/o Multi-Agent Collaboration: Use a single agent, which incorporates all prompts from our four proposed agents as a unified workflow, with our proposed scaling and reward strategies retained. (3) w/o Agent-Wise Hybrid Scaling: Retain the multi-agent collaborative framework and apply the mixed reward strategy directly, without test-time scaling. (4) w/o Mixed Reward Modeling: Retain the multi-agent collaborative framework and agent-wise hybrid scaling, without reward signals. As shown in Table 2, the multi-agent collaborative approach contributes the most, outperforming the single-agent system by 8.6%. Notably, consolidating all functions into a single model results in suboptimal performance, which is even worse than the base model, highlighting the necessity of a multi-agent collaborative framework. Both the mixed reward modeling and agent-wise hybrid scaling strategies boost average accuracy by at least 3%. Specifically, the latter yields more significant gains on complex tasks compared to average ones, with especially notable improvements in reasoning tasks.
Analysis of Multi-Agent Collaboration
To assess the utility of each agent and their collaboration, we perform additional experiments on different agent combinations, which are listed in Table 3. The combination of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$ results in a 3.9% average increase over the base model, while adding $\mathcal{A}_{judg}$ further boosts performance by 5.5%. $\mathcal{A}_{ans}$ extraly improves the scores by 0.9%, with this gain expanding to 1.5% for long visual context. Additionally, as in Figure 4(a), we compare our judgment agent strategy with the other two counterparts. Our proposed strategy outperforms the others by at least 2.6% on average, while requiring an average of 0.3 fewer corrections. Moreover, our method achieves optimal correction results when the maximum number of corrections $N_{c}$ is set to 3, whereas the alternatives require $N_{c}$ to be set to 5, highlighting its greater accuracy and efficiency. It could also be observed that when setting a higher $N_{c}$, the performance will not improve continuously, primarily because excessive corrections may confuse the agent and obscure correct answers.
MMLong TabBen MVision Average
No Reward 38.3 54.2 36.7 71.4
Agent-Specific Reward 42.8 56.3 40.5 72.7
Global Reward 34.1 51.6 35.1 70.2
@@
Agent-Wise Hybrid 43.7 57.2 41.8 74.8
Scaling & Mixed Reward
Table 4: Results of different reward modeling and test-time scaling strategies.
@@
Figure 4: (a) Line graph shows the impact of various maximum numbers of corrections, with dashed and dotted lines denoting average values across all and three selected benchmarks, respectively. The bar charts show the average judgment numbers when the maximum is set to 3. (b) The line graphs represent the impacts of $N_{p}$ and $N_{e}$.
Analysis of Reward modeling and Test-Time Scaling
@@
We perform experiments on our proposed mixed reward modeling and agent-wise hybrid test-time scaling. For the former, we separately use the agent-wise reward and global reward for comparison. For the latter, we benchmark our method with the other four effective test-time scaling strategies. Here, we select budget forcing (Muennighoff et al. 2025) as the internal scaling. Although the improvement from global reward is relatively limited, it avoids the selfishness of agents that only use agent-specific reward, yielding a further 1.3% increase in average scores. Our strategy, which designs tailored scaling forms for each agent according to its functions and specificity, outperforms all four existing strategies and improves the average scores by at least 1.8%. As demonstrated in Figure 4(b), higher values of $N_{p}$ and $N_{e}$ increase the likelihood of finding the correct answers, which leads to more attempts at plans and execution steps.
Conclusion

### v2

@@ 2.2 Agent-Wise Adaptive Test-Time Scaling
For visual document understanding and reasoning, the key challenge is not visual documents themselves, but the procedural reasoning that links global question decomposition, plan formation, information extraction, and mistake judgment across multiple steps. Monolithic scaling enlarges model capacity uniformly, yet fails to accommodate this structured reasoning process. We therefore propose an agent-wise adaptive scaling strategy during inference, dynamically allocating additional computing resources to different functional entities and enabling on-demand allocation during the test time. Existing test-time scaling methods fall into four main categories: parallel scaling, sequential scaling, hybrid scaling, and internal scaling. However, these strategies are originally designed for single models or agents, overlooking the different division of labor among agents and achieving suboptimal performance when applied to multi-agent systems. Accordingly, we propose an agent-wise adaptive scaling strategy tailored to multi-agent architectures, which significantly enhances both agent-specific capabilities and collaborative performance.
@@
For the first three agents with various functions, we customize different test-time scaling strategies to their unique characteristics: (1) [Uncaptioned image] Planning agent: In our design of $\mathcal{A}_{plan}$, it naturally provides relevant sample plans as references for formulating multiple execution plans for the original task. We therefore prompt the generation of $N_{p}$ relevant plans independently, yielding $N_{p}$ parallel paths per question. This establishes a basis for scaling subsequent agents within document processing workflows, increasing the probability that at least one reasoning path will produce an accurate answer aligned with document semantics. (2) [Uncaptioned image] Execution agent: Given that the execution process is divided by step in $\mathcal{A}_{exe}$, we treat each step as an evaluation node. For each node, the agent produces $\mathit{N_{e}}$ candidate executions, which are scored using a pretrained reward model. The top-scoring candidate is selected as the base node for subsequent steps, with all others rejected. (3) [Uncaptioned image] Judgment agent: To judge the correctness of the execution plans and processes, $\mathcal{A}_{judg}$ requires logical analysis and reasoning to comprehensively and accurately detect mistakes to avoid the error snowballing in document contexts. We therefore adopt the budget forcing scaling method [muennighoff2025s1] for this agent, which enforces a minimum number of thinking tokens. When token usage falls below the budget, the agent is encouraged to generate additional thinking tokens, thereby promoting accurate judgments. For the answer agent, whose core function lies in synthesizing prior information and generating the final response, test-time scaling confers merely marginal improvements.
2.3 Mixed Reward Modeling
According to the latest research [setlur2025scaling, zhang2025survey], adding RL strategy, e.g., reward modeling, to optimize test-time scaling is significantly superior to fine-tuning or training-free methodologies. To optimize our proposed procedural scaling by elevating task-agent-centric and collaborative performance, we design a mixed reward strategy that synthesizes agent-specific rewards and global outcome-driven reward signals. Given that the four agents in the framework have distinct functions and reward signal preferences, we first incorporate agent-tailored rewards:
$r_{plan/exe}$ $=\mathcal{R}_{prm}\left({s}_{i}/{e}_{i}\mid\mathcal{Q},\mathcal{D},\mathcal{P}/\mathcal{E}\right),$ (6)
$r_{judg/ans}$ $=\mathcal{R}_{orm}\left(\mathcal{J}/\mathcal{O}\mid\mathcal{Q}\right),$ (7)
where are $\mathcal{R}_{prm}$ and $\mathcal{R}_{orm}$ are multi-modal process reward and outcome reward models, respectively. The former yields step-wise process rewards, providing instant feedback for each step and hierarchical rewards for $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, and the latter generates single reward signal for each output of $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$. Furthermore, we apply a global reward based on the final selected path of the four agents:
$r_{global}=\mathcal{R}_{orm}\left(\left\{\mathcal{P},\mathcal{E},\mathcal{J},\mathcal{O}\right\}\mid\mathcal{Q},\mathcal{D}\right),$ (8)
which reinforces incentives for correct paths and moderately mitigates agent self-interest. Since individual agents tend to optimize their own local rewards, the integration of a global objective effectively alleviates such self-serving tendencies. Please refer to Appendix 7.3 for details.
Table 1: Comparisons of our MACT and other counterparts on 15 benchmarks. The table cells with green, blue, and gray colors denote the best, the second best, and the third best values. Results with - indicate exceeding the maximum context token limit or not being equipped with a specific ability. $\triangle,\Box,\Diamond$ correspond to the base models of three variants of MACT for comparisons.
Document Non-Document Avg.
@@ 2.3 Mixed Reward Modeling
[…]
3 Experiments
More details and descriptions about the training pipelines, datasets, evaluations and prompts for each agent are provided in Appendix 8.1, 8.2, 8.3, and 8.4, respectively.
3.1 Training Pipeline
@@
Since the functions and optimization goals of each agent are independent and diverse, we design a two-stage SFT and RL pipeline for MACT. To understand visual inputs and better accomplish their labors, we choose VLMs for the $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, while LLMs for the $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$. For all the four agents, we initiate with pretrained models, and we select three groups of small-parameter base models for different variants of MACT: (1) Qwen2.5-VL series based [bai2025qwen2, qwen2025qwen25technicalreport]: Qwen2.5-VL-7B-Instruct and Qwen2.5-7B/3B-Instruct; (2) MiMo-VL series based [coreteam2025mimovl, xia2025mimo]: MiMo-VL-7B-SFT and MiMo-7B-SFT; (3) InternVL3 series based [zhu2025internvl3]: InternVL3-9B/8B/2B-Instruct. In the first SFT stage, we initially train a tuned 11B/7B/7B VLM on the selected document-based or non-document-based datasets, mixing data with or without CoT. It aims to enhance their visual understanding and reasoning abilities, providing more robust models for future multi-agent collaboration. Next, we fine-tune an 8B/7B/7B LLM on judgment labels generated via GPT-4o [hurst2024gpt] and rule-based verifications. Finally, we fine-tune another 3B/3B/7B LLM on the outputs from preceding agents and ground-truths to enhance its summary ability. The fine-tuned VLM are applied as $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$ respectively in the subsequent stage, while the two LLMs function as the $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$. In the second RL stage, we generate reward signals based on pretrained reward models and optimize our model via GRPO [shao2024deepseekmath]. Process reward model VisualPRM [wang2025visualprm] is utilized for the step-by-step reward signals of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, while Skywork-VL-Reward [wang2025skywork] is used for the rewards generation of $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$.
3.2 Benchmarks
@@
To comprehensively and objectively evaluate our model’s document-based capabilities, we selected 15 datasets encompassing four real-world document categories and two non-document general categories. The four document types are comprised of: (1) Text-based: DocVQA [mathew2021docvqa], DUDE [van2023document], SlideVQA [tanaka2023slidevqa], MMLongBench-Doc [ma2025mmlongbench]; (2) Webpage-based: VisualMRC [tanaka2021visualmrc], InfographicVQA [mathew2022infographicvqa]; (3) Chart-based: ChartQA [masry2022chartqa], CharXiv [wang2024charxiv]; (4) Table-based: TableVQA-Bench [kim2024tablevqa], and TableBench [wu2025tablebench]. To validate our document-centric paradigm without sacrificing capabilities in general domains, two non-document types are involved: (1) General: ScienceQA [saikh2022scienceqa], RealWorldQA [2024XAIrealworld]; (2) Mathematical: MathVista [lu2024mathvista], Math-Vision [wang2024measuring], and MathVerse [zhang2024mathverse]. We adhere to the original training and testing splits and use the testing splits as evaluation benchmarks. Each instance in these datasets consists of a visual input sequence and a question, spanning difficulty levels from easy to hard and encompassing various question types for comprehensive evaluation.
3.3 Evaluations
@@
For other reproducible models, we utilize LMMs-Eval [zhang2024lmms] for fair comparisons on both natively supported and our registered benchmarks. For most benchmarks, we employ GPT-4o [hurst2024gpt] as a judge model to evaluate the correctness of the generated answers based on LMMs-Eval. For the remaining benchmarks, we utilize their original evaluation metrics, e.g., ANLS and F1, as detailed in Tab. 8.
4 Results and Discussions
4.1 Main Results
@@
We evaluate MACT across 15 benchmarks: 10 document-based benchmarks to assess its document understanding performance; 2 general benchmarks and 3 mathematical benchmarks are designed to verify that it retains full capabilities in non-document-centric and reasoning task settings. For comparison, we select state-of-the-art (SOTA) methods, encompassing both generalist and specialist models, and categorize them by parameter size.
@@
Benchmark Results. As shown in Tab. 1, the MACT-MiMo-VL-Series-28B variant delivers the best average performance, followed by MACT-InternVL3-Series-28B and MACT-Qwen2.5-VL-Series-24B, which rank second and third, respectively. Despite having fewer than 30B parameters, MACT models outperform all comparative methods with under 100B parameters, as well as closed-source models. Notably, MACT-MiMo-VL-Series-28B achieves significant improvements of 5.6% and 5.9% over the top-performing open-source and closed-source models in terms of the average scores. Additionally, the three variants top 13 of the 15 benchmarks, with MACT-MiMo-VL-Series-28B leading on seven. Notably, in MMLongBench-Doc, which features the longest visual context, and across the three mathematical reasoning benchmarks, MACT-MiMo-VL-Series-28B outperforms the second-highest scorer by 7.1%, 10.6%, 5.9%, and 8.7%, respectively. These results highlight that document-based tasks, far from being monolithic constructs, are better suited to procedural scaling than to monolithic models with larger parameters, supporting flexible adaptation to functional entities and computational resource allocation.
@@
Comparisons with Base Models. As shown in Fig. 1, radar charts compare three variants of our proposed MACT framework with their base models and larger models from the same series. It is observed that MACT models significantly outperform their base models by 10.3%, 9.9%, and 11.5% on average across the 15 benchmarks. Besides, compared to their corresponding larger monolithic models, i.e., Qwen2.5-VL-72B-Instruct and InternVL3-78B-Instruct, our MACT-Qwen2.5-VL-Series-24B and MACT-InternVL3-Series-28B achieve average performance gains of at least 6.6% and 3.7%, respectively, with more pronounced improvements observed on long-context and reasoning tasks. These results demonstrate that our method, equipped with procedural scaling paradigm, is superior to monolithic paradigm in the document-based tasks and also more general domains.
4.2 Additional Quantitative Analysis
@@
To further demonstrate the effectiveness of our proposed procedural scaling paradigm, we perform ablation experiments and analyses of scaling components. Experiments here are conducted using MACT-Qwen2.5-VL-Series-24B, and the training pipelines and training data are consistent with the main experiment. We select three lowest-performing benchmarks, distinguished by long visual contexts and reasoning-intensive demands, and the average performance across all benchmarks, aiming to quantify their impacts on both specialized capabilities.
Table 2: Results of ablations on multi-agent collaboration, agent-wise adaptive test-time scaling, and mixed reward modeling.
MMLong TabBen MVision Avg.
Monolithic 32.5 50.8 32.4 66.2
w/o Multi-Agent Collaboration 24.7 44.4 26.0 58.6
w/o Agent-Wise Adaptive Scaling 34.9 50.8 34.5 71.1
w/o Mixed Reward Modeling 38.3 54.2 36.7 71.4
@@
Table 3: Results of different combinations of agents.
$\mathcal{A}_{plan}$ + $\mathcal{A}_{exe}$ $\mathcal{A}_{judg}$ $\mathcal{A}_{ans}$ MMLong TabBen MVision Avg.
@@
Ablation Studies. The ablation settings are as follows: (1) Monolithic: Use a monolithic model to directly execute tasks and output answers, with particular scaling and reward modeling retained. (2) w/o multi-agent collaboration: Use a single agent, which incorporates all prompts from our four proposed agents as a unified workflow, with our proposed scaling and reward strategies retained. (3) w/o agent-wise adaptive test-time scaling: Retain the multi-agent collaborative framework and apply the mixed reward strategy directly, without test-time scaling. (4) w/o mixed reward modeling: Retain the multi-agent collaborative framework and agent-wise adaptive test-time scaling, without reward signals. As reported in Tab. 2, our document-centric procedural scaling paradigm, outperforming the monolithic system by 8.6%. Notably, the integration of all functionalities into a single model results in suboptimal performance, even outperforming the base model negatively, which indicates the inherent limitation of monolithic paradigms in document understanding and reasoning. Both the agent-wise adaptive test-time scaling strategy and mixed reward modeling contribute to an average accuracy increase of at least 3%. Specifically, the former yields more remarkable gains on reasoning tasks in documents than average ones, with particularly striking improvements that underscore how adaptive scaling reduces cognitive overload.
Table 4: Results of different setting of test-time scaling strategies.
MMLong TabBen MVision Avg.
@@
Agent-Wise Adaptive 43.7 57.2 41.8 74.8
Table 5: Results of different setting of reward modeling.
MMLong TabBen MVision Avg.
No Reward 38.3 54.2 36.7 71.4
Agent-Specific Reward 42.8 56.3 40.5 72.7
Global Reward 34.1 51.6 35.1 70.2
Mixed Reward 43.7 57.2 41.8 74.8
Analysis of Multi-Agent Collaboration. To assess the utility of each agent and the effectiveness their procedural collaboration, we perform additional experiments on different agent combinations, which are demonstrated in Tab. 3. The combination of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, as a basic procedural multi-agent system, results in a 3.9% average increase over the base model, while adding $\mathcal{A}_{judg}$ further boosts performance by 5.5%. $\mathcal{A}_{ans}$ additionally improves the scores by 0.9% on the average. Additionally, as in Fig. 4(a), we compare our judgment agent strategy with the other two counterparts, i.e., internal correction and agent for both judgment and correction. Our proposed strategy outperforms the others by at least 2.6% on average, while requiring an average of 0.3 fewer corrections. Moreover, our method achieves optimal correction results when the maximum number of corrections $N_{c}$ is set to 3, whereas the alternatives require $N_{c}$ to be set to 5, highlighting its greater accuracy and efficiency. It could also be observed that when setting a higher $N_{c}$, the performance will not improve continuously, primarily because excessive corrections may confuse the agent and obscure correct answers.
@@
Figure 4: (a) Line graph shows the impact of various maximum numbers of corrections, with solid and dashed lines denoting average values across all and three selected benchmarks, respectively. The bar charts show the average judgment numbers when the maximum is set to 3. (b) The line graphs represent the impacts of the number of generated plans $\mathit{N_{p}}$ and candidate executions $\mathit{N_{e}}$.
@@
Analysis of Agent-Wise Adaptive Test-Time Scaling and Mixed Reward modeling. We perform experiments on our proposed agent-wise adaptive test-time scaling and mixed reward modeling. For the former, we benchmark our method with the other four effective test-time scaling strategies, as listed in Tab. 4. Here, we select budget forcing [muennighoff2025s1] as the internal scaling. Our agent-wise adaptive scaling strategy based on the computational resources and cognitive load of each agent, outperforms all four existing strategies and improves the average scores by at least 1.8%. As shown in Fig. 4(b), higher values of $\mathit{N_{p}}$ and $\mathit{N_{e}}$ increase the likelihood of finding the correct answers, which leads to more attempts at plans and execution steps. For the latter, we separately use the agent-wise reward and global reward for comparison, as reported in Tab. 5. Although the improvement from global reward is relatively limited, it avoids the selfishness of agents that only use agent-specific reward, yielding a further 1.3% increase in average scores.
5 Conclusion